# The Marked Edge Walk: A Novel MCMC Algorithm for Sampling of Graph Partitions

### v1

@@ 1 Introduction
Despite great attention to the spanning tree distribution, there is still an open problem of sampling from any distribution. Metropolized Forest Recombination (MFR) [17] attempts to do exactly that, utilizing the powerful tools of the Metropolis-Hastings algorithm to target any desired invariant distribution. To achieve this, the authors ‘lift’ to an expanded state space of spanning forests that allows for easier calculation of the forward and reverse transition probabilities. However, this algorithm showed poor mixing behavior when sampling from distributions that were not proportional to the spanning tree distribution.
@@ 4 Target Distribution
$\tau(\xi)=\Pi_{i=1}^{d}t(\xi_{i})\cdot\Pi_{\xi_{i}\sim\xi_{j}}|\partial(\xi_{i},\xi_{j})|$ (8)
where $t(\xi_{i})$ counts the number of spanning trees in component $\xi_{i}$, and $|\partial(\xi_{i},\xi_{j})|$ denotes the number of edges between adjacent components $\xi_{i}$ and $\xi_{j}$.
@@
This choice of target distribution allows us to target any desired distribution over $P_{d}(G)$ while sampling in the lifted space $\mathcal{X}$. One reason this is so important is that, especially on large dual graphs, the spanning tree count of partitions is so astronomically large that it dwarfs other constraints that we might try to build into an energy function. For example, in Figure 2, we show two possible redistricting plans of Texas’s congressional districts. The map on the left is drawn from the uniform distribution on $\mathcal{X}$, corresponding to the spanning-tree/border-length distribution, and the map on the right is the 2020 enacted plan. The sampled plan has a degeneracy factor of $\tau(\xi_{\text{sampled}})\approx 10^{5160}$, and the enacted plan has a degeneracy factor of $\tau(\xi_{\text{enacted}})\approx 10^{4694}$. Thus, when sampling from the spanning-tree/border-length distribution, the sampled plan is about $10^{466}$ times more likely to be sampled than the enacted plan. To put the number $10^{466}$ into perspective: there are an estimated $10^{80}$ atoms in the observable universe. So, the sampled plan is more likely than the enacted plan by a factor that is astronomically larger than the number of atoms in the universe.
@@ 5 Results
We evaluate the efficacy of MEW through numerical experiments on three real-world graphs with varying energy functions $J$. First, we target the uniform distribution over partitions of the dual graph of Cheshire County, an example that permits comparison to an enumerated baseline. Second, we target the spanning-tree/border-length distribution on the dual graph of New Hampshire, and since ReCom draws independent samples from this distribution when $d=2$, we can use it as a baseline. Third, we target competitiveness, compactness, and multivariate distributions over the dual graph of New Hampshire. Finally, we target a multivariate distribution on the dual graph of Texas to validate MEW’s scalability to large dual graphs with many districts. Our results demonstrate strong evidence of mixing in all cases.
@@ 5.2 The Spanning Tree Distribution
We investigate the spanning-tree/border-length distribution as a target distribution for two primary reasons. Since much attention has been devoted to its study, researchers may want to use MEW to target the distribution or related distributions. Secondly, since many algorithms are available to target the distribution, we can use it as another baseline to assess convergence. To this end, we employ ReCom, which, in the $d=2$ case, draws independent samples from the spanning-tree/border-length distribution; each partition has probability proportional to the product of the spanning tree product and the boundary length. Notice that this is exactly the uniform distribution on $\mathcal{X}$, thus we set:
@@
As we can see in Figure 4, MEW successfully samples from this distribution. The average pairwise Kolmogorov-Smirnov (KS) distance between the chains decays to nearly zero (0.0124 after 2.5 million steps), and the average KS distance from the chains to the baseline also decays to nearly zero (0.0188 after 2.5 million steps). The figure also reveals a limitation of sampling from the spanning-tree/border-length distribution and related distributions. When sampling less compact partitions to target realistic plans (the enacted plan has 72 cut edges) or to investigate the effects of highly noncompact districts (targeting plans with 100 or more cut edges), we would see long mixing times and high rejection rates due to the distribution’s naturally compact structure.
@@ 6 Conclusion and Discussion
In this paper we introduce the marked edge walk (MEW), a novel MCMC algorithm for sampling from the space of balanced graph partitions under a tuneable target distribution. The procedure combines a step of the cycle basis walk on spanning trees with a marked edge transition, and this balance of global and local steps allows for good exploration of the state space while maintaining easily calculable transition probabilities. Unlike previous approaches such as the Metropolized flip walk or Metropolized Forest ReCom (MFR), MEW mixes well under target distributions unrelated to the spanning tree distribution.
When targeting the uniform distribution on the dual graph of Cheshire County, the spanning-tree/border-length distribution and a competitiveness distribution on the dual graph of New Hampshire, the chain exhibits fast mixing evidenced by rapidly decaying pairwise KS distance and KS distance to the target distribution.

### v2

@@ 1 Introduction
Despite great attention to the spanning tree distribution, there is still an open problem of sampling from any distribution. Metropolized Forest Recombination (MFR) [17] attempts to do exactly that, utilizing the powerful tools of the Metropolis-Hastings algorithm to target any desired invariant distribution. To achieve this, the authors ‘lift’ to an expanded state space of spanning forests that allows for easier calculation of the forward and reverse transition probabilities. However, it is still an open question what distributions can effectively be sampled using this method. Recently, a preprint was released that introduces a novel walk on the lifted space that appears to reduce the reliance on the spanning tree distribution [18].
@@ 4 Target Distribution
$\tau(\xi)=\Pi_{i=1}^{d}t(\xi_{i})\cdot t(Q)$ (8)
where $t(\xi_{i})$ counts the number of spanning trees in component $\xi_{i}$, and $Q=G\setminus\xi$ is the quotient graph.
@@
This choice of target distribution allows us to target any desired distribution over $P_{d}(G)$ while sampling in the lifted space $\mathcal{X}$. One reason this is so important is that, especially on large dual graphs, the spanning tree count of partitions is so astronomically large that it dwarfs other constraints that we might try to build into an energy function. For example, in Figure 2, we show two possible redistricting plans of Texas’s congressional districts. The map on the left is drawn from the uniform distribution on $\mathcal{X}$, corresponding to the spanning tree distribution, and the map on the right is the 2020 enacted plan. The sampled plan has a degeneracy factor of $\tau(\xi_{\text{sampled}})\approx 10^{5160}$, and the enacted plan has a degeneracy factor of $\tau(\xi_{\text{enacted}})\approx 10^{4694}$. Thus, when sampling from the spanning tree distribution, the sampled plan is about $10^{466}$ times more likely to be sampled than the enacted plan. To put the number $10^{466}$ into perspective: there are an estimated $10^{80}$ atoms in the observable universe. So, the sampled plan is more likely than the enacted plan by a factor that is astronomically larger than the number of atoms in the universe.
@@ 5 Results
We evaluate the efficacy of MEW through numerical experiments on three real-world graphs with varying energy functions $J$. First, we target the uniform distribution over partitions of the dual graph of Cheshire County, an example that permits comparison to an enumerated baseline. Second, we target the spanning tree distribution on the dual graph of New Hampshire, and since ReCom draws independent samples from this distribution when $d=2$, we can use it as a baseline. Third, we target competitiveness, compactness, and multivariate distributions over the dual graph of New Hampshire. Finally, we target a multivariate distribution on the dual graph of Texas to validate MEW’s scalability to large dual graphs with many districts. Our results demonstrate strong evidence of mixing in all cases.
@@ 5.2 The Spanning Tree Distribution
We investigate the spanning tree distribution as a target distribution for two primary reasons. Since much attention has been devoted to its study, researchers may want to use MEW to target the distribution or related distributions. Secondly, since many algorithms are available to target the distribution, we can use it as another baseline to assess convergence. To this end, we employ ReCom, which, in the $d=2$ case, draws independent samples from the spanning tree distribution; each partition has probability proportional to the product of the spanning tree product of each part and the spanning tree product of the quotient graph. Notice that this is exactly the uniform distribution on $\mathcal{X}$, thus we set:
@@
As we can see in Figure 4, MEW successfully samples from this distribution. The average pairwise Kolmogorov-Smirnov (KS) distance between the chains decays to nearly zero (0.0124 after 2.5 million steps), and the average KS distance from the chains to the baseline also decays to nearly zero (0.0188 after 2.5 million steps). The figure also reveals a limitation of sampling from the spanning tree distribution and related distributions. When sampling less compact partitions to target realistic plans (the enacted plan has 72 cut edges) or to investigate the effects of highly noncompact districts (targeting plans with 100 or more cut edges), we would see long mixing times and high rejection rates due to the distribution’s naturally compact structure.
@@ 6 Conclusion and Discussion
In this paper, we introduce the marked edge walk (MEW), a novel MCMC algorithm for sampling from the space of balanced graph partitions under a tuneable target distribution. The procedure combines a step of the cycle basis walk on spanning trees with a marked edge transition, and this balance of global and local steps allows for good exploration of the state space while maintaining easily calculable transition probabilities. Unlike previous approaches such as the Metropolized flip walk or Metropolized Forest ReCom (MFR), MEW mixes well under target distributions unrelated to the spanning tree distribution.
When targeting the uniform distribution on the dual graph of Cheshire County, the spanning tree distribution, and a competitiveness distribution on the dual graph of New Hampshire, the chain exhibits fast mixing evidenced by rapidly decaying pairwise KS distance and KS distance to the target distribution.
@@ 7 Data and Code Availability
8 Acknowledgments
This work was funded by the Dartmouth Fellowship to Atticus McWhorter and by the Alfred P. Sloan Foundation (G-2021-16778) and the National Science Foundation (DMS-1928930) to Daryl DeFord.